# CRPS-Based Targeted Sequential Design with Application in Chemical Space

### v1

@@ 3.3 SUR selection criteria
$=\mathbb{E}_{n,\boldsymbol{x}}\left[\int\mathrm{CRPS}_{\nu}((\xi_{n+1}(\boldsymbol{x^{\prime}}))\mathbb{P}_{\mathcal{X}}(\mathrm{d}\boldsymbol{x^{\prime}})\right]$
$=\int\mathbb{E}_{n,\boldsymbol{x}}\left[\mathrm{CRPS}_{\nu}((\xi_{n+1}(\boldsymbol{x^{\prime}}))\right]\mathbb{P}_{\mathcal{X}}(\mathrm{d}\boldsymbol{x^{\prime}}).$
This means we want to select $\boldsymbol{x_{n+1}}$ such that it minimizes the expected integrated future $\mathrm{CRPS}_{\nu}(\boldsymbol{\xi}_{n+1})$. We continue by deriving expressions for the integrand, keeping $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$ fixed. In doing so, we consider the marginal Gaussian $\xi_{n+1}(\boldsymbol{x^{\prime}})\sim\mathcal{N}(m_{n+1}(\boldsymbol{x^{\prime}}),k_{n+1}(\boldsymbol{x^{\prime}},\boldsymbol{x^{\prime}}))$. While the updated variance $k_{n+1}(\boldsymbol{x^{\prime}},\boldsymbol{x^{\prime}})$ only depends on $\boldsymbol{x_{n+1}}=\boldsymbol{x}$, the updated mean $m_{n+1}(\boldsymbol{x^{\prime}})$ also depends on the noise-contaminated corresponding value $z_{n+1}=f(\boldsymbol{x_{n+1}})+\epsilon_{n+1}$ (Eq. 20). When we consider adding $\boldsymbol{x_{n+1}}=\boldsymbol{x}$, we neither can nor want to evaluate all candidate points in advance, so we do not yet know $z_{n+1}$. However, using the expectation $\mathbb{E}_{n,\boldsymbol{x}}$ with respect to $\boldsymbol{\xi_{n}}$, we can write $Z_{n+1}=m_{n}(\boldsymbol{x})+\sqrt{k_{n}(\boldsymbol{x},\boldsymbol{x})+\tau^{2}}V$ with $V\sim\mathcal{N}(0,1)$ such that
@@
We consider the conditional GP $\boldsymbol{\xi}_{n+1}\sim\mathrm{GP}(m_{n+1},k_{n+1})$ and its predecessor $\boldsymbol{\xi_{n}}\sim\mathrm{GP}(m_{n},k_{n})$. For fixed $\boldsymbol{x},\boldsymbol{x^{\prime}}\in\mathcal{X}$ and using $\sigma^{2}_{n+1}=k_{n+1}(\boldsymbol{x^{\prime}},\boldsymbol{x^{\prime}})$, $\mu_{n}=m_{n}(\boldsymbol{x^{\prime}})$, $\alpha_{n}=\alpha_{n}(\boldsymbol{x^{\prime}})$, we obtain for $\nu_{1}$:
$\mathbb{E}_{n,\boldsymbol{x}}\left[\mathrm{CRPS}_{\nu_{1}}(\xi_{n+1}(\boldsymbol{x^{\prime}}))\right]=\sigma_{n+1}\mathbb{E}\left[\left(N-\max\left(\widetilde{N},\frac{t-\mu_{n}-\alpha_{n}V}{\sigma_{n+1}}\}\right)\right)_{+}\right],$
with $N,\widetilde{N},V\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. For $\nu_{2}$, one derives,
$\mathbb{E}_{n,\boldsymbol{x}}\left[\mathrm{CRPS}_{\nu_{2}}(\xi_{n+1}(\boldsymbol{x^{\prime}}))\right]$ $=\Phi\left(\begin{pmatrix}0\\ 0\end{pmatrix};\begin{pmatrix}\mu_{n}-t\\ t-\mu_{n}\end{pmatrix},\begin{pmatrix}\sigma_{n+1}^{2}+\alpha_{n}^{2}+\sigma_{\nu}^{2}&-\alpha_{n}^{2}-\sigma_{\nu}^{2}\\ -\alpha_{n}^{2}-\sigma_{\nu}^{2}&\sigma_{n+1}^{2}+\alpha_{n}^{2}+\sigma_{\nu}^{2}\end{pmatrix}\right).$
@@
Appendix A.3.3. ∎
Using these formulas for the integrand, the SUR criteria $J_{n}(\boldsymbol{x})$ can be calculated by integrating over $\boldsymbol{x^{\prime}}\in\mathcal{X}$. In a discrete molecular space, this corresponds to taking the mean of the molecules. We refer to the SUR criteria, which are based on integration, with $\mathrm{ICRPS}_{\nu_{1}}$ and $\mathrm{ICRPS}_{\nu_{2}}$.
@@ 4.2 Tanimoto Kernel
Following Thawani et al., (2020), we use fingerprints as input representations. Given the nature of these high-dimensional binary inputs, we not only consider popular traditional kernels (as presented in Appendix A.2), but also kernels that are specialized for such non-continuous input spaces. For a comprehensive comparison of kernels, including alternative molecular representations, we refer to Griffiths et al., (2023).
@@
$k_{\mathrm{TAN}}(\boldsymbol{x},\boldsymbol{x^{\prime}}):=\left\{\begin{array}[]{ll}\sigma_{k}^{2}&\text{ if }\langle\boldsymbol{x},\boldsymbol{x}\rangle=\langle\boldsymbol{x^{\prime}},\boldsymbol{x^{\prime}}\rangle=0,\\ \sigma_{k}^{2}\frac{\langle\boldsymbol{x},\boldsymbol{x^{\prime}}\rangle}{||\boldsymbol{x}||^{2}+||\boldsymbol{x^{\prime}}||^{2}-\langle\boldsymbol{x},\boldsymbol{x^{\prime}}\rangle}&\text{ else, }\end{array}\right.$ (11)
@@ 4.3 Implementation and configuration
The main code for this study can be found on GitHub
(https://github.com/LeaFrie/GP_sequentialCRPS)
and the original Photoswitch dataset is publicly available
(see e.g., https://paperswithcode.com/dataset/photoswitch).
For implementing the GP models, we utilized the R-package kergp (Deville et al.,, 2021). This is a package for GP regression where some pre-defined kernels are available, and it is also possible to define customized kernels through a formula mechanism. The latter was employed in the case of the Tanimoto kernel. We consider the variance term of the kernel (such as the integral scale if there is one) as hyperparameters, estimated simultaneously using maximum likelihood estimation with numerical optimization via gradient based methods. In the test case concerning the original data with the noise being unknown, we consider its variance as an additional hyperparameter. For the sequential data acquisition, we implemented the different criteria manually. For the integral criteria, we calculate the integrals by averaging over the candidate set. Furthermore, for $\mathrm{ICRPS}_{\nu_{1}}$, we employ Monte Carlo sampling of the standard-normally distributed random variables.
The synthetic test case is constructed using a Tanimoto kernel, so we naturally also use it in the later numerical experiments. For the choice of kernel used on the original dataset, we compared the performance of the Tanimoto, Gaussian and Exponential kernels using $10,20,30\%$ of the datasets for training (Appendix A.2). Given the consistently good performance of the Tanimoto kernel, even with small training datasets, we continue focusing on this kernel, also for the original dataset. For the threshold $t$ specifying the excursion set, we employ the empirical 0.8-quantile of the distribution of the numerical response of interest to ensure that we have a reasonable number of molecules with the desired property in the dataset.
@@ 4.4 Performance evaluation
To prevent a scenario where all excursion set molecules are sequentially added to the training set, leaving no examples in the test set, we first reserve a subset of molecules for validation in each trial. This leads to a split of the data into three sets: the initial training set with points $\boldsymbol{x_{1}},\dots,\boldsymbol{x_{n}}$, the candidate training set $\boldsymbol{x_{n+1}},\dots,\boldsymbol{x_{m}}$ and the validation set with points $\boldsymbol{x_{m+1}},\dots,\boldsymbol{x_{N}}$. We specify the number of validation points as $M=N-m=100$ in both test cases. Eventually, the predictive performance of a model is evaluated on the validation set by considering the mean metric of the validation points.
In the test case which we constructed such that the ground truth is known, we consider the following metrics: Naturally, we use the CRPS-based evaluation metrics (Section 3), wherein we compare the latent GP $\Phi_{\mu,\sigma^{2}}$ with mean $\mu=m_{n}(\boldsymbol{x})$ and variance $\sigma^{2}=k_{n}(\boldsymbol{x,\boldsymbol{x}})$ with the ground truth $y=f(\boldsymbol{x})$. To avoid confusion, we denote the threshold-weighted CRPS evaluation metrics as $\overline{\mathrm{CRPS}}_{\nu_{1}}$ and $\overline{\mathrm{CRPS}}_{\nu_{2}}$, indicating that we consider the mean value over the validation molecules. Furthermore, we consider the precision and sensitivity of the model given by,
@@
where we use the numbers of true positives (TP; $|\hat{\Gamma}\cap\Gamma|$), false positives (FP; $|\hat{\Gamma}\cap\Gamma^{c}|$), true negatives (TN; $|\hat{\Gamma}^{c}\cap\Gamma^{c}|$) and false negatives (FN; $|\hat{\Gamma}^{c}\cap\Gamma|$), with $|A|$ denoting the cardinality of set $A$ and $A^{c}$ representing its complement. Additionally, we consider two versions of the Root Mean Squared Error (RMSE), which should not be confused with the acquisition criterion based on the Mean Squared Error (kriging variance). The first version exclusively considers molecules actually lying in the excursion set,
[…]
@@
[…]
@@ 5 Results
In this section, we evaluate the performance of the new acquisition strategies and compare it against state-of-the-art methods and a random approach, where the additional training molecules are selected without an acquisition function. To make reading easier amidst the collection of threshold-weighted CRPS terms, we recall that $\overline{\mathrm{CRPS}}_{\nu_{1}}$ and $\overline{\mathrm{CRPS}}_{\nu_{2}}$ are used to denote the evaluation metrics (Section 3.1), $\mathrm{CRPS}{\nu_{1}}$ and $\mathrm{CRPS}{\nu_{2}}$ refer to the pointwise acquisition criteria (Section 3.2) and $\mathrm{ICRPS}{\nu_{1}}$ and $\mathrm{ICRPS}{\nu_{2}}$ to the SUR criteria (Section 3.3).
@@ 5.1 Synthetic dataset
We start with the synthetic test case. Figure 3 shows the evolution of the evaluation metrics with $n=30$ initial molecules and adding $\widetilde{n}=50$ more molecules. We depict the mean over 60 repetitions, which are all started with another random initial training set. We observe that, with few exceptions, the acquisition criteria consistently yield better results compared to the random approach. In terms of (a) $\overline{\mathrm{CRPS}}_{\nu_{1}}$ and (d) $\overline{\mathrm{CRPS}}_{\nu_{2}}$ ($\sigma_{\nu}=33$), the corresponding $\mathrm{ICRPS}{\nu_{1}}$ and $\mathrm{ICRPS}{\nu_{2}}$ ($\sigma_{\nu}=33$) perform the best. However, the difference between the pointwise and SUR criteria is more pronounced for $\overline{\mathrm{CRPS}}_{\nu_{2}}$, while for $\overline{\mathrm{CRPS}}_{\nu_{1}}$, they perform comparably after 15 steps. Furthermore, after 20 steps, the pointwise criterion $\mathrm{CRPS}{\nu_{1}}$ performs best with respect to (b) $\mathrm{RMSE}_{p}$ and (e) precision, but worst for (f) sensitivity and (d) $\overline{\mathrm{CRPS}}_{\nu_{2}}$. Regarding (e) $\mathrm{RMSE}_{\Gamma}$, both $\mathrm{ICRPS}{\nu_{1}}$ and $\mathrm{ICRPS}{\nu_{2}}$ perform about equally well and better than the pointwise criteria, while for (f) sensitivity, $\mathrm{CRPS}{\nu_{2}}$ followed by $\mathrm{ICRPS}{\nu_{2}}$ achieve the highest values.
@@
Figure 3: The sequential performance on the synthetic dataset with $M=100$ validation molecules, $n=30$ and $\widetilde{n}=50$; the mean over 60 repetitions is depicted.
@@
To better understand the process, Figure 4 shows an example of an (a) initial model and the evolving performance for $\widetilde{n}=20$ using (b) $\mathrm{CRPS}{\nu_{1}}$ (c) and $\mathrm{CRPS}{\nu_{2}}$. The dark red dots represent the original (noisy) training set, the blue crosses indicate the predicted mean for the $M=100$ validation molecules, the light red circles indicate the added molecules, and the gray dots show the ground truth $f$. It can be clearly seen that while $\mathrm{CRPS}{\nu_{1}}$ excels in accurately predicting the responses of the TP, it is less sensitive and still has more FN compared to $\mathrm{CRPS}{\nu_{2}}$. We observe that $\mathrm{CRPS}{\nu_{1}}$ primarily adds molecules from the excursion set, while $\mathrm{CRPS}{\nu_{2}}$ focuses on molecules near the threshold, as expected based on the two weighting measures.
@@
Figure 4: One illustrative example with the synthetic dataset, starting from (a) $n=30$ and adding $\widetilde{n}=20$ molecules with (b) $\mathrm{CRPS}_{\nu_{1}}$ and (c) $\mathrm{CRPS}_{\nu_{2}}$. The darkred dots show the original (noisy) training set, the blue crosses the predictive mean for the $M=100$ validation molecules, the lightred circles the added molecules and the grey dots in the back ground the ground truth $f$.
@@
Figure 5 shows the performance of $\mathrm{CRPS}_{\nu_{2}}$ and $\mathrm{ICRPS}_{\nu_{2}}$ as a function of $\sigma_{\nu}$. For both criteria, the traditional CRPS in Figure 5(a) decreases as expected with increasing $\sigma_{\nu}$ (because then the criteria focus more and more on the whole domain, as the CRPS does). The pointwise $\mathrm{CRPS}_{\nu_{2}}$ shows a clear improvement in (b) sensitivity and a deterioration in (c) precision with increasing $\sigma_{\nu}$ (after an increase in the first point for the later). The same is also present for the SUR criterion with respect to precision, less so for sensitivity. In all previous and following results we use $\sigma_{\nu}=33$, which is half the standard deviation of the responses of the whole dataset.
@@
Figure 5: Performance of $\mathrm{CRPS}_{\nu_{2}}$ and $\mathrm{ICRPS}_{\nu_{2}}$ using different $\sigma_{\nu}$ for the synthetic dataset; $M=100$ validation molecules, mean over 60 repetitions, $n=30$ and $\widetilde{n}=25$.
@@
In Figure 6, we compare the performance of the new acquisition criteria against the state-of-the-art methods. The box and whiskers plots display the distribution of the metrics obtained in 60 repetitions, highlighting the range, median and quartiles. Also in comparison with the alternative methods, the pointwise $\mathrm{CRPS}_{\nu_{1}}$ performs best with respect to the (a) $\overline{\mathrm{CRPS}}_{\nu_{1}}$ (reduction of the median compared to the random strategy by about 25 %), (b) $\mathrm{RMSE}_{p}$ (minus 40 % compared to random) and (c) precision (10 % higher than random). However, $\mathrm{ICRPS}_{\nu_{1}}$ leads to a similar performance with respect to the median in $\overline{\mathrm{CRPS}}_{\nu_{1}}$ and precision. For (d) $\overline{\mathrm{CRPS}}_{\nu_{2}}$, the performance of $\mathrm{ICRPS}_{\nu_{2}}$ is comparable to that of TIMSE, which makes sense since TIMSE also relies on a Gaussian weight around the threshold. Both reduce the median value of the random strategy by about 15 %. Regarding (e) $\mathrm{RMSE}_{\Gamma}$, the best median value is observed for TMSE (minus 20 %). The median of the (f) sensitivity is the highest for TMSE (15 % higher than for random), followed by $\mathrm{CRPS}_{\nu_{2}}$ and TIMSE.
@@
Figure 6: Performance comparison between $\mathrm{CRPS}_{\nu_{1}}$, $\mathrm{CRPS}_{\nu_{2}}$, $\mathrm{ICRPS}_{\nu_{1}}$ and $\mathrm{ICRPS}_{\nu_{2}}$ against alternative selection criteria using the synthetic dataset and $M=100$ validation molecules (60 repetitions, $n=30$ and $\widetilde{n}=25$). The box and whiskers plots depict the median, quartiles, two whiskers and all ’outlying’ points. The whiskers extend from the quartiles to the smallest / highest value at most 1.5 $\times$ the interquartile range away from the quartile, beyond which points are considered outliers.

### v2

@@ 3.3 SUR selection criteria
$=\mathbb{E}_{n,\boldsymbol{x}}\left[\int\mathrm{CRPS}_{\gamma}(\xi_{n+1}(\boldsymbol{x^{\prime}}))\mathbb{P}_{\mathcal{X}}(\mathrm{d}\boldsymbol{x^{\prime}})\right]$
$=\int\mathbb{E}_{n,\boldsymbol{x}}\left[\mathrm{CRPS}_{\gamma}(\xi_{n+1}(\boldsymbol{x^{\prime}}))\right]\mathbb{P}_{\mathcal{X}}(\mathrm{d}\boldsymbol{x^{\prime}}).$
This means we want to select $\boldsymbol{x_{n+1}}$ such that it minimizes the expected integrated future $\mathrm{CRPS}_{\gamma}(\boldsymbol{\xi}_{n+1})$. We continue by deriving expressions for the integrand, keeping $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$ fixed. In doing so, we consider the marginal Gaussian $\xi_{n+1}(\boldsymbol{x^{\prime}})\sim\mathcal{N}(m_{n+1}(\boldsymbol{x^{\prime}}),k_{n+1}(\boldsymbol{x^{\prime}},\boldsymbol{x^{\prime}}))$. While the updated variance $k_{n+1}(\boldsymbol{x^{\prime}},\boldsymbol{x^{\prime}})$ only depends on $\boldsymbol{x_{n+1}}=\boldsymbol{x}$, the updated mean $m_{n+1}(\boldsymbol{x^{\prime}})$ also depends on the noise-contaminated corresponding value $z_{n+1}=f(\boldsymbol{x_{n+1}})+\epsilon_{n+1}$. When we consider adding $\boldsymbol{x_{n+1}}=\boldsymbol{x}$, we neither can nor want to evaluate all candidate points in advance, so we do not yet know $z_{n+1}$. However, using the expectation $\mathbb{E}_{n,\boldsymbol{x}}$ with respect to $\boldsymbol{\xi_{n}}$, we can write $Z_{n+1}=m_{n}(\boldsymbol{x})+\sqrt{k_{n}(\boldsymbol{x},\boldsymbol{x})+\tau^{2}}V$ with $V\sim\mathcal{N}(0,1)$ such that
@@
We consider the conditional GP $\boldsymbol{\xi}_{n+1}\sim\mathrm{GP}(m_{n+1},k_{n+1})$ and its predecessor $\boldsymbol{\xi_{n}}\sim\mathrm{GP}(m_{n},k_{n})$. For fixed $\boldsymbol{x},\boldsymbol{x^{\prime}}\in\mathcal{X}$ and using $\sigma^{2}_{n+1}=k_{n+1}(\boldsymbol{x^{\prime}},\boldsymbol{x^{\prime}})$, $\mu_{n}=m_{n}(\boldsymbol{x^{\prime}})$, $\alpha_{n}=\alpha_{n}(\boldsymbol{x^{\prime}})$, we obtain for $\gamma_{1}$:
$\mathbb{E}_{n,\boldsymbol{x}}\left[\mathrm{CRPS}_{\gamma_{1}}(\xi_{n+1}(\boldsymbol{x^{\prime}}))\right]=\sigma_{n+1}\mathbb{E}\left[\left(N-\max\left(\widetilde{N},\frac{t-\mu_{n}-\alpha_{n}V}{\sigma_{n+1}}\}\right)\right)_{+}\right],$
with $N,\widetilde{N},V\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. For $\gamma_{2}$, one derives,
$\mathbb{E}_{n,\boldsymbol{x}}\left[\mathrm{CRPS}_{\gamma_{2}}(\xi_{n+1}(\boldsymbol{x^{\prime}}))\right]$ $=\Phi\left(\begin{pmatrix}0\\ 0\end{pmatrix};\begin{pmatrix}\mu_{n}-t\\ t-\mu_{n}\end{pmatrix},\begin{pmatrix}\sigma_{n+1}^{2}+\alpha_{n}^{2}+\sigma_{\gamma}^{2}&-\alpha_{n}^{2}-\sigma_{\gamma}^{2}\\ -\alpha_{n}^{2}-\sigma_{\gamma}^{2}&\sigma_{n+1}^{2}+\alpha_{n}^{2}+\sigma_{\gamma}^{2}\end{pmatrix}\right)$.
@@
Supplementary material (Part C.3). ∎
Using these formulas for the integrand, the SUR criteria $J_{n}(\boldsymbol{x})$ can be calculated by integrating over $\boldsymbol{x^{\prime}}\in\mathcal{X}$. In a discrete molecular space, this corresponds to taking the mean of the molecules. We refer to the SUR criteria, which are based on integration, with $\mathrm{ICRPS}_{\gamma_{1}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$.
To ensure the theoretical soundness of these SUR criteria, we establish in Part D of the supplementary material that $\mathrm{ICRPS}_{\gamma_{1}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$ are consistent in the sense of Bect et al., (2019). Specifically, we show that the corresponding uncertainty measures $\mathrm{CRPS}_{\gamma_{1}}$ and $\mathrm{CRPS}_{\gamma_{2}}$ converge almost surely to zero as the number of observations increases, thereby ensuring the asymptotic correctness of the sequential design.
@@ 4.2 Tanimoto Kernel
Following Thawani et al., (2020), we use fingerprints as input representations. Given the nature of these high-dimensional binary inputs, we not only consider popular traditional kernels (as presented in the supplementary material), but also kernels that are specialized for such non-continuous input spaces. For a comprehensive comparison of kernels, including alternative molecular representations, we refer to Griffiths et al., (2023).
@@
[…]
@@ 4.3 Implementation and configuration
The code for this study can be found on GitHub, and the original Photoswitch dataset is publicly available (see e.g., Photoswitch dataset). For implementing the GP models, we utilized the R-package kergp (Deville et al.,, 2021). This is a package for GP regression where some pre-defined kernels are available, and it is also possible to define customized kernels through a formula mechanism. The latter was employed in the case of the Tanimoto kernel. We consider the variance term of the kernel (such as the integral scale if there is one) as hyperparameters, estimated simultaneously using maximum likelihood estimation with numerical optimization via gradient based methods. In the test case concerning the original data with the noise being unknown, we consider its variance as an additional hyperparameter. For the sequential data acquisition, we implemented the different criteria manually. For the integral criteria, we calculate the integrals by averaging over the candidate set. Furthermore, for $\mathrm{ICRPS}_{\gamma_{1}}$, we employ Monte Carlo sampling of the standard-normally distributed random variables.
The synthetic test case is constructed using a Tanimoto kernel, so we naturally also use it in the later numerical experiments. For the choice of kernel used on the original dataset, we compared the performance of the Tanimoto, Gaussian and Exponential kernels using $10,20,30\%$ of the datasets for training (supplementary material, Part B). Given the consistently good performance of the Tanimoto kernel, even with small training datasets, we continue focusing on this kernel, also for the original dataset. For the threshold $t$ specifying the excursion set, we employ the empirical 0.8-quantile of the distribution of the numerical response of interest to ensure that we have a reasonable number of molecules with the desired property in the dataset.
@@ 4.4 Performance evaluation
To prevent a scenario where all excursion set molecules are sequentially added to the training set, leaving no examples in the test set, we first reserve a subset of molecules for validation in each trial. This leads to a split of the data into three sets: the initial training set with points $\boldsymbol{x_{1}},\dots,\boldsymbol{x_{n}}$, the candidate training set $\boldsymbol{x_{n+1}},\dots,\boldsymbol{x_{m}}$ and the validation set with points $\boldsymbol{x_{m+1}},\dots,\boldsymbol{x_{N}}$. We specify the number of validation points as M=Nm=100M=N-m=100 in both test cases. Eventually, the predictive performance of a model is evaluated on the validation set by considering the mean metric of the validation points.
In the test case which we constructed such that the ground truth is known, we consider the following metrics: Naturally, we use the CRPS-based evaluation metrics (Section 3), wherein we compare the latent GP $\Phi_{\mu,\sigma^{2}}$ with mean $\mu=m_{n}(\boldsymbol{x})$ and variance $\sigma^{2}=k_{n}(\boldsymbol{x,\boldsymbol{x}})$ with the ground truth $y=f(\boldsymbol{x})$. To avoid confusion, we denote the evaluation metrics as $\overline{\mathrm{CRPS}}$, $\overline{\mathrm{CRPS}}_{\gamma_{1}}$ and $\overline{\mathrm{CRPS}}_{\gamma_{2}}$, indicating that we consider the mean value over the validation molecules. Furthermore, we consider the precision and sensitivity of the model given by,
@@
where we use the numbers of true positives (TP; $|\hat{\Gamma}\cap\Gamma|$), false positives (FP; $|\hat{\Gamma}\cap\Gamma^{c}|$), true negatives (TN; $|\hat{\Gamma}^{c}\cap\Gamma^{c}|$) and false negatives (FN; $|\hat{\Gamma}^{c}\cap\Gamma|$), with $|A|$ denoting the cardinality of set $A$ and $A^{c}$ representing its complement. Additionally, we consider the Root Mean Squared Error (RMSE), which should not be confused with the acquisition criterion based on the Mean Squared Error (kriging variance). Besides the basic version, we also employ two versions targeting the excursion scenario. The first version exclusively considers molecules actually lying in the excursion set,
$\mathrm{RMSE_{\Gamma}}=\sqrt{\frac{1}{M_{\Gamma}}\sum_{i=m+1}^{N}(m_{n}(\boldsymbol{x_{i}})-f(\boldsymbol{x_{i}}))^{2}\,\mathbbm{1}\{\boldsymbol{x_{i}}\in\Gamma\}},$ (13)
@@
$\mathrm{RMSE_{p}}=\sqrt{\frac{1}{M_{p}}\sum_{i=m+1}^{N}(m_{n}(\boldsymbol{x_{i}})-f(\boldsymbol{x_{i}}))^{2}\,\mathbbm{1}\{\boldsymbol{x_{i}}\in\hat{\Gamma}\}}$, (14)
@@ 4.5 Alternative selection criteria
For classification problems, margin-based active learning approaches rely on the distance of inputs to the decision boundary (Tong and Koller,, 2001; Brinker,, 2003; Balcan et al.,, 2007; Ducoffe and Precioso,, 2018). To refine the current model, these methods query unlabeled inputs that lie close to the decision boundary, as they are expected to provide the most information for improving the classifier. In the case of binary classification, the entropy-based acquisition criterion can be viewed as related to margin-based active learning approaches, with the exceedance probability reflecting the closeness of an input to the decision boundary.
@@ 5 Results
In this section, we evaluate the performance of the new acquisition strategies and compare it against state-of-the-art methods and a random approach, where the additional training molecules are selected without an acquisition function. To make reading easier amidst the collection of threshold-weighted CRPS terms, we recall that $\overline{\mathrm{CRPS}}_{\gamma_{1}}$ and $\overline{\mathrm{CRPS}}_{\gamma_{2}}$ are used to denote the evaluation metrics (Section 3.1), $\mathrm{CRPS}_{\gamma_{1}}$ and $\mathrm{CRPS}_{\gamma_{2}}$ refer to the pointwise acquisition criteria (Section 3.2) and $\mathrm{ICRPS}_{\gamma_{1}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$ to the SUR criteria (Section 3.3).
@@ 5.1 Synthetic dataset
We start with the synthetic test case. Figure 3 shows the evolution of the evaluation metrics with $n=30$ initial molecules and adding $\widetilde{n}=50$ more molecules. We depict the mean over 100 repetitions, which are all started with another random initial training set and based on a different validation set. We observe that on the excursion set–specific evaluation metrics (a-f), the acquisition criteria consistently outperform the random approach (with few exceptions). In terms of (a) $\overline{\mathrm{CRPS}}_{\gamma_{1}}$, both corresponding $\mathrm{CRPS}_{\gamma_{1}}$ and $\mathrm{ICRPS}_{\gamma_{1}}$ and perform equally well. For (d) $\overline{\mathrm{CRPS}}_{\gamma_{2}}$ ($\sigma_{\gamma}=33$), the SUR criterion $\mathrm{ICRPS}_{\gamma_{2}}$ performs noticeably better that pointwise $\mathrm{CRPS}_{\gamma_{2}}$. Furthermore, after 20 steps, the pointwise criterion $\mathrm{CRPS}_{\gamma_{1}}$ performs best with respect to (b) $\mathrm{RMSE}_{p}$ and (e) precision, but worst for (f) sensitivity and (d) $\overline{\mathrm{CRPS}}_{\gamma_{2}}$. Regarding (e) $\mathrm{RMSE}_{\Gamma}$, $\mathrm{ICRPS}_{\gamma_{2}}$ performs best, while for (f) sensitivity, $\mathrm{CRPS}_{\gamma_{2}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$ achieve the highest values. Finally, $\mathrm{ICRPS}_{\gamma_{2}}$ performs best with respect to both ‘basic’ (g) $\overline{\mathrm{CRPS}}$ and (h) RMSE.
@@
Figure 3: The sequential performance on the synthetic dataset with $M=100$ validation molecules, $n=30$ and $\widetilde{n}=50$; the mean over 100 repetitions is depicted and for better visual clarity, only every second step is shown.
@@
To better understand the process, Figure 4 shows an example of an (a) initial model and the evolving performance for $\widetilde{n}=25$ using (b) $\mathrm{CRPS}_{\gamma_{1}}$ (c) and $\mathrm{CRPS}_{\gamma_{2}}$. The filled triangles represent the original (noisy) training set, the crosses indicate the predicted mean for the $M=100$ validation molecules, the filled boxes indicate the added molecules, and the dots in the background show the ground truth $f$. It can be clearly seen that while $\mathrm{CRPS}_{\gamma_{1}}$ excels in accurately predicting the responses of the TP, it is less sensitive and still has more FN compared to $\mathrm{CRPS}_{\gamma_{2}}$. We observe that $\mathrm{CRPS}_{\gamma_{1}}$ primarily adds molecules from the excursion set, while $\mathrm{CRPS}_{\gamma_{2}}$ focuses on molecules near the threshold, as expected based on the two weighting measures.
@@
Figure 4: One illustrative example with the synthetic dataset, starting from (a) $n=30$ and adding $\widetilde{n}=25$ molecules with (b) $\mathrm{CRPS}_{\gamma_{1}}$ and (c) $\mathrm{CRPS}_{\gamma_{2}}$. The filled triangles show the original (noisy) training set, the crosses the predictive mean for the $M=100$ validation molecules, the filled boxes the added molecules and the dots in the back ground the ground truth $f$.
@@
Figure 5 shows the performance of $\mathrm{CRPS}_{\gamma_{2}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$ as a function of $\sigma_{\gamma}$. For both criteria, the traditional CRPS in Figure 5(a) decreases as expected with increasing $\sigma_{\gamma}$ (because then the criteria focus more and more on the whole domain, as the CRPS does). Both criteria show a clear deterioration in (c) precision with increasing $\sigma_{\gamma}$. For (b) sensitivity, there is a slight improvement for the pointwise $\mathrm{CRPS}_{\gamma_{2}}$ with increasing $\sigma_{\gamma}$, but no systematic effect for $\mathrm{ICRPS}_{\gamma_{2}}$. In all previous and following results we use $\sigma_{\gamma}=33$, which is half the standard deviation of the responses of the whole dataset.
@@
Figure 5: Performance of $\mathrm{CRPS}_{\gamma_{2}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$ using different $\sigma_{\gamma}$ for the synthetic dataset; $M=100$ validation molecules, mean over 50 repetitions, $n=30$ and $\widetilde{n}=25$.
@@
In Figure 6, we compare the performance of the new acquisition criteria against the state-of-the-art methods. The box and whiskers plots display the distribution of the metrics obtained in 100 repetitions, highlighting the range, median and quartiles. Also in comparison with the alternative methods, the pointwise $\mathrm{CRPS}_{\gamma_{1}}$ performs good with respect to the (a) $\overline{\mathrm{CRPS}}_{\gamma_{1}}$ (reduction of the median compared to the random strategy by about 30 %), (b) $\mathrm{RMSE}_{p}$ (minus 40 % compared to random) and (c) precision (10 % higher than random). However, $\mathrm{ICRPS}_{\gamma_{1}}$ leads to a similar performance with respect to the median in $\overline{\mathrm{CRPS}}_{\gamma_{1}}$ and the entropy criteria performs slightly better for precision. For (d) $\overline{\mathrm{CRPS}}_{\gamma_{2}}$, the good performance of $\mathrm{ICRPS}_{\gamma_{2}}$ is comparable to that of TIMSE, which makes sense since TIMSE also relies on a Gaussian weight around the threshold. Both reduce the median value of the random strategy by about 15 %. Regarding (e) $\mathrm{RMSE}_{\Gamma}$, the best median value is observed for TMSE (minus 16 %), closely followed by $\mathrm{CRPS}_{\gamma_{2}}$, $\mathrm{ICRPS}_{\gamma_{2}}$ and TIMSE. The median of the (f) sensitivity is the highest for $\mathrm{CRPS}_{\gamma_{2}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$ (15 % higher than for random), followed by TIMSE and IBV. For reasons of completeness, we also show the results in ‘basic’ (g) $\overline{\mathrm{CRPS}}$ and (h) RMSE. For both metrics, the medians of $\mathrm{ICRPS}_{\gamma_{2}}$ and TIMSE are the lowest, reducing the value of the random strategy by 10 %.
@@
Figure 6: Performance comparison between $\mathrm{CRPS}_{\gamma_{1}}$, $\mathrm{CRPS}_{\gamma_{2}}$, $\mathrm{ICRPS}_{\gamma_{1}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$ against alternative selection criteria using the synthetic dataset and $M=100$ validation molecules (100 repetitions, $n=30$ and $\widetilde{n}=25$). The box and whiskers plots depict the median, quartiles, two whiskers and all ’outlying’ points. The whiskers extend from the quartiles to the smallest / highest value at most 1.5 $\times$ the interquartile range away from the quartile, beyond which points are considered outliers.